### v1

# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

Abstract
While the progression of Large Language Models (LLMs) has notably propelled financial analysis, their application has largely been confined to singular language realms, leaving untapped the potential of bilingual Chinese-English capacity. To bridge this chasm, we introduce ICE-PIXIU [^1] https://github.com/The-FinAI/PIXIU, seamlessly amalgamating the ICE-INTENT model and ICE-FLARE benchmark for bilingual financial analysis. ICE-PIXIU uniquely integrates a spectrum of Chinese tasks, alongside translated and original English datasets, enriching the breadth and depth of bilingual financial modeling. It provides unrestricted access to diverse model variants, a substantial compilation of diverse cross-lingual and multi-modal instruction data, and an evaluation benchmark with expert annotations, comprising 10 NLP tasks, 20 bilingual specific tasks, totaling 1,185k datasets. Our thorough evaluation emphasizes the advantages of incorporating these bilingual datasets, especially in translation tasks and utilizing original English data, enhancing both linguistic flexibility and analytical acuity in financial contexts. Notably, ICE-INTENT distinguishes itself by showcasing significant enhancements over conventional LLMs and existing financial LLMs in bilingual milieus, underscoring the profound impact of robust bilingual data on the accuracy and efficacy of financial NLP.
1 Introduction
Inspired by the success of general-domain large language models (LLMs), the exploration of LLMs in the financial sector is gaining momentum Shah et al. (2022); Lu et al. (2023); Xie et al. (2023b, 2024, a). Starting with the broad capabilities of models like GPT-3.5 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), the focus has shifted towards more specialized models such as BloombergGPT (Wu et al., 2023), which tackles the complex nuances of financial language and concepts. Subsequently, models like FinGPT (Yang et al., 2023b), InvestLM (Yang et al., 2023c) and Typhoon (Pipatanakul et al., 2023) emerged, concentrating on English and Thai financial tasks. Notably, FinMA  (Xie et al., 2023b) addresses the challenges of instruction data modality within this domain, marking a significant advancement.
Beyond the efforts in English-centric models, Chinese financial LLMs (FinLLMs) have also made notable advances. XuanYuan2.0 (Zhang, 2023) exemplifies this trend, focusing on the intricacies of Chinese financial language. This has been followed by models such as DISC-FinLLM (Chen et al., 2023) and CFGPT (Li et al., 2023a), which further the capabilities in Chinese financial analysis and reasoning. Recently, PanGu-π (Wang et al., 2023) enhances a pretrained model for better performance in Chinese financial examinations. Moreover, certain business-oriented LLMs like TigerBot and LightGPT remain closed-source (Wang, 2023).
Model Size Language Instruction Evaluation Open Source Modality Release Date
zh en zh en dof Model Data Only Text
BloombegGPT (Wu et al., 2023) 50B en 0 5 0 5 0 $usym2717 \usym$2714 03/30/23
InvestLM (Yang et al., 2023c) 65B en 0 9 0 9 0 $usym2717 \usym$2714 09/14/23
FinGPT (Wu et al., 2023) 7/13B en 0 6 0 6 0 $usym2717 \usym$2714 11/10/23
XuanYuan2.0 (Zhang, 2023) 176B zh - 0 0 1 0 $usym2717 \usym$2714 05/19/23
FinMA (Xie et al., 2023b) 7/13B en 0 8 0 8 0 $usym2714 \usym$2717 06/01/23
[…]
ICE-INTENT 7B zh, en 12 6 12 6 3 $usym2714 \usym$2717 12/10/23
[…]
However, there is a significant challenge in their application and assessment in bilingual contexts, especially regarding model development, instruction dataset diversity, and evaluation methodologies in English and Chinese. As Table 1 illustrates, most FinLLMs are tailored exclusively for either English or Chinese, which underscores a significant shortfall in models proficient in both languages.
This issue is further compounded in the domain of instruction datasets. While there are tasks encompassing classification, extraction, prediction and reasoning, these datasets, particularly for Chinese, show a notable lack of diversity in each task category (Chen et al., 2023; Li et al., 2023a). Furthermore, the original English datasets and their subsequent translations into Chinese, are often overlooked in the development of comprehensive bilingual instruction datasets (Xie et al., 2023b).
Correspondingly, the evaluation methodologies for these models tend to mirror these linguistic limitations (Zhang et al., 2023b; Lei et al., 2023; Zhang et al., 2023c), focusing mainly on monolingual assessments and failing to comprehensively evaluate bilingual proficiency. There is a critical need for more diverse Chinese instruction datasets in both classification, extraction, reasoning and prediction, alongside a greater focus on the translation and utilization of original English datasets into Chinese.
To address this issue, we introduce ICE-PIXIU, a comprehensive framework that features ICE-INTENT and ICE-FLARE, the first cross-lingual bilingual financial LLM and evaluation benchmark, respectively. This framework is characterized by several key attributes:
  • Bilingual Capability Proficiency: ICE-INTENT, a component of ICE-PIXIU, demonstrates outstanding bilingual proficiency in English and Chinese, essential for global financial data processing and analysis.
  • Diverse Chinese Datasets: ICE-PIXIU addresses gaps in Chinese financial NLP by incorporating a variety of Chinese classification, extraction, reasoning and prediction NLP tasks, enhancing training and performance.
  • Professionally Annotated Prompts: ICE-PIXIU offers a a diverse, high-quality, expert-annotated set of prompts applied with similar fine-tuning instructions to enhance understanding of financial tasks.
  • Incorporation of Translation and English Data: The framework expands its capabilities by including translation tasks and English datasets, bolstering its bilingual training and application.
  • Cross-Lingual Evaluation with ICE-FLARE: ICE-PIXIU introduces ICE-FLARE, a rigorous cross-lingual evaluation benchmark, ensuring consistent model performance in diverse linguistic contexts.
  • Open-Source Contribution: With its open-access approach, ICE-PIXIU offers its resources to the research community, fostering collaborative advancement in financial NLP.
In the creation of ICE-INTENT, we meticulously gathered 40 datasets, which consisted of 1,185,076 raw data, 603,940 instruction fine-tuning data, and 95,091 evaluation data, strategically covering a wide range of financial tasks. This ensemble, as detailed in Table 1, comprises 18 specific task on Chinese-English bilingual instruction datasets. This diverse array not only underscores our model’s commitment to linguistic breadth but also to task-specific depth, ensuring ICE-INTENT’s proficiency across various financial scenarios  [^2] Detailed examples are shown in Appendix Figure 7-8. Complementing this, ICE-FLARE, our cross-lingual evaluation benchmark, incorporates 21 evaluation specific tasks—12 in Chinese and 9 in English. Particularly, we innovatively introduced data-out-of-field specific task into ICE-FLARE. This robust framework highlights the unique strength of our approach in providing comprehensive cross-lingual consistency and generalization assessments, crucial for bilingual FinLLMs.
In our evaluation conducted using ICE-FLARE, we observed distinct performance patterns of our financial LLM ICE-INTENT in comparison with advanced general-domain LLMs and existing FinLLMs. The results firstly emphasized the model’s enhanced capabilities in cross-lingual tasks, particularly in handling complex and lengthy text, underscoring the critical role of bilingual data integration in financial language modeling. Subsequently, the model exhibited proficiency in translation tasks, highlighting its adeptness at managing linguistic variations crucial for global financial contexts. Finally, our model outperformed general LLMs in tasks specifically fine-tuned for the financial domain but faced challenges in less familiar tasks within ICE-FLARE. This delineates the specialized nature and unique challenges of the financial sector, reinforcing the need for tailored FinLLMs.
2 Related Work
Financial Dataset Resource: Financial pre-trained language models (FinPLMs) previously focused on single-task data but now most FinLLMs prioritize specific-task data collection. For instance, BloombergGPT (Wu et al., 2023) integrates general and financial text for training large 500B LLMs. FinLLMs like FinGPT (Yang et al., 2023b; Zhang et al., 2023a) curate English instruction datasets for classification, extraction, and generation NLP tasks. PIXIU  (Xie et al., 2023b) provide open-source, multi-task and multi-modal data covering diverse financial scenarios. BBT-Fin (Lu et al., 2023) Corpus aggregates Chinese financial data including corporate announcements, research reports, social media, and more, while XuanYuan (轩辕) 2.0 (Zhang, 2023) offers a hybrid-tuning dataset. Cornucopia’s (Yu, 2023) 26M instruction dataset covers application areas like insurance and stocks. Most fine-tuning datasets are tailored for monolingual tasks, limiting coverage and open sourcing.
Financial Language Models: FinPLMs and FinLLMs build upon backbone models like BERT (Devlin et al., 2018) and LLaMa (Touvron et al., 2023) with large-scale financial texts. FinBERT (Araci, 2019) was the first FinPLM with open released corpus, later improved FinBERT (Yang et al., 2020) incorporating larger communication datasets. Models like FLANG Shah et al. (2022) combine BERT and ELECTRA (Clark et al., 2020). Chinese-specific PLMs like Mengzi (Zhang et al., 2021) and FinT5 (Lu et al., 2023) address unique financial tasks. BloombergGPT introduce the outstanding FinLLM, along with successors like FinGPT (Yang et al., 2023b), DISC-FinLLM (Chen et al., 2023), CFGPT (Li et al., 2023a), and FinMA (Xie et al., 2023b) excelling in specific data collection and supervised fine-tuning using backbone models Baichuan (Yang et al., 2023a) and InternLM (Team, 2023). However, monolingual FinLLMs struggle to provide unified cross-lingual finance services.
Financial Evaluation Benchmarks: The first benchmark FLUE (Shah et al., 2022) and its successor FLARE (Xie et al., 2023b) evaluate classification, boundary detection, and question answering abilities in English domain. FLARE also evaluates time-series reasoning like forecasting stock price movements. BBT-CFLEB (Lu et al., 2023) is the first Chinese benchmark with 5 financial NLP tasks, while FinCUGE  (Chen et al., 2023) focuses on 9 financial NLP tasks on social media and short news. FinEval (Zhang et al., 2023b) assesses LLMs through multiple-choice questions on finance, economy, and accounting. CFBenchmark  (Lei et al., 2023) evaluates text processing abilities in recognition, classification, and generation, encompassing totaling 8 tasks, and CGCE (Zhang et al., 2023c) includes 200 general and 150 specific professional diverse QA tasks. However, most benchmarks lack comprehensive real-world application support, posing challenges for assessing bilingual FinLLMs.
3 Method
The overall framework of ICE-PIXIU, displayed in Figure 1, offers tailored services to various financial groups like banks, funds, and insurance firms. It offers professional solutions like event analysis, Q&A, stock forecasting, exams, credit ratings. ICE-PIXIU’s robust analytical prowess stems from meticulously curated 5 data types—Data Label from Classification (DLC), Data Label from Extraction (DLE), Data Text by Translation (DTT), Data Text with English (DTE), and Data Out of Field (DOF). These categories support various NLP tasks, encompassing classification (CLS), extraction (EXT), reasoning (REA), prediction (PRE), generation (GEN) and Translation (TRA), across 15 bilingual specific financial tasks in Chinese and English. Moreover, ICE-PIXIU not only boasts bilingual proficiency but also its broad support for multi-modal features, such as tables in financial reports and time series data in stock predictions.
Figure 1: An overview of ICE-PIXIU, its unique data types, financial tasks, and sources in the bilingual domain of Chinese and English, providing services for different user groups in various financial scenarios.
3.1 Raw Data Construction
In this section, we introduce our bilingual Chinese-English raw data, designed for instruction tuning and evaluation benchmark. Compared with similar works,  (Chen et al., 2023) focuses on CLS and REA financial NLP task,  (Lei et al., 2023; Li et al., 2023a) emphasizes CLS and EXT financial NLP task. Our raw data comprises 40 different datasets covering almost all financial NLP tasks, including CLS, EXT, REA, PRE, GEN, and TRA. Unlike commonly self-instruct generation data methods  (Wang et al., 2022), our data is sourced from open datasets in real-world financial scenarios. The sunburst chart illustrating the distribution of ICE-PIXIU across different levels is shown in Figure 2. As depicted, FinED and FinSM tasks boast datasets on a larger scale. Furthermore, Chinese data surpasses English data in terms of proportions in data types, NLP tasks, and specific tasks, particularly showcasing a more diverse distribution in specific financial tasks, covering a wider range of financial application scenarios. This indicates that compared to the English-focused PIXIU (Xie et al., 2023b), ICE-PIXIU demonstrates superior understanding of Chinese financial tasks. More details information of our raw data are in Appendix Table 6. According to statistics, the distribution percentages for all data types are: DLE (59.38%), DLC (20.08%), DTT (14.24%), DTE (5.96%), and DOF (0.34%), while the distribution percentages for Chinese and English NLP tasks are: ZH-CLS (18.83%), ZH-PRE (1.25%), ZH-EXT (56.64%), ZH-GEN (2.73%), ZH-TRA (14.24%), EN-CLS (1.61%), EN-PRE (3.31%), EN-EXT (0.05%), EN-REA (0.99%), and EN-DOF (0.34%).
Figure 2: A sunburst chart depicting the distribution of ICE-PIXIU across language proficiency, data types, financial NLP tasks, specific financial tasks, and datasets.
3.2 Bilingual Financial Datasets
The ICE framework’s proficiency in bilingualism is benefited from the following Chinese financial data, English financial data, and Translation data.
3.2.1 Chinese Financial Datasets
We mainly focus on Chinese language proficiency. Thus, for Chinese, we provide two data types distinguished by their source of the labels: DLC and DLE, which encompass classification (ZH-CLS), extraction (ZH-EXT), prediction (ZH-PRE) and generation (ZH-GEN) task in financial NLP.
For ZH-CLS, our dataset covers essential tasks such as sentiment analysis (FinSA), semantic matching (FinSM), news classification (FinNC), negative judgment (FinNJ), answer selection (FinAS) and relationship extraction (FinRE). 1) Sentiment Analysis. It is the process of analyzing and labeling the sentiment expressed in financial texts as "Negative(消极)","Neutral(中性)" and "Negative(消极)" (Sohangir et al., 2018). We used the FE (Lu et al., 2023) dataset sourced from Stockbar and Xueqiu websites. We also utilized the StockB   [^3] https://huggingface.co/datasets/kuroneko5943/stock11 dataset, which focuses on the finance in among data from ten domains. 2) Semantic Matching. It determines if financial texts or sentences are semantically similar ("是(YES)" or "否(NO)"), using two datasets: the Bank Question Corpus (BQC)  (Chen et al., 2018) from Chinese bank customer logs and the Ant Financial Question Matching Corpus (AFQMC)  (Xu et al., 2020) from the Alipay competition, both utilizing binary classification for match identification. 3) Negative Judgment. It focuses on identifying finance and economics-related negativity, distinct from general financial sentiment analysis. Leveraging the FinNSP dataset (Lu et al., 2023) sourced from social media, it uses "是(YES)" to label data containing negative information and "否(NO)" for those without. 4) News Classification. It is to categorize news into 2 or multiple-class label, such as "股票市场(stock market)", “外汇(foreign exchange)", "宏观经济(macroeconomics), and more. We use the FinNL  (Lu et al., 2023) dataset, sourced from financial news websites like ina Finance, Tencent Finance, Phoenix Finance, 36Kr, and Huxiu. It includes binary (NL) and multi-class (NL2) labeled data. 5) Answer Selection. FinEval  (Zhang et al., 2023b) and and AntGroup evaluates financial reasoning by selecting the correct option ("A", "B", "C", "D) in a financial context. The dataset   [^4] https://github.com/alipay/financial_evaluation_dataset includes 4,661 questions covering finance, economy, accounting, and certifications. We’ve compiled data for ten key finance subjects as FinEvalF. 6) Relationship Extraction. It aims to provide the types of connections, dependencies, and associations between entities. We gathered the FinRE dataset (Lu et al., 2023) and named it RE, which includes financial news articles along with head-tail entity pairs, covering 44 relationship categories (e.g., "合作(cooperation)", "持股(shareholding)", etc.), including the "未知(unknown)" relation.
For ZH-EXT, our dataset includes question answering (FinQA), named entity recognition (FinER), and event detection (FinED). 1) Question Answering. It aims to extract accurate answers from the given text to answer questions related to an event mentioned in the text. We have collected a dataset named QA to differentiate from the English dataset FinQA (Lu et al., 2023) with the same name. 2) Named Entity Recognition(NER). It aims to identify entity names and types labeled as individuals ("个人(PER)"), organizations ("组织(ORG)"), or locations ("地点(LOC)") in financial text. We utilize the the comprehensive NER dataset, referred to as CNER in (Jia et al., 2020). 3) Event Detection. It involves recognizing and comprehending specific incidents beyond industry categorization. We collect four different financial challenge datasets   [^5] https://www.biendata.xyz/competition/{ccks_2019_4,ccks_2020_4_1,ccks_2021_task6_2,ccks2022_eventext} from China Conference on Knowledge Graph and Semantic Computing (CCKS), named as 19CCKS, 20CCKS, 21CCKS and 22CCKS, respectively. Unlike similar work (Lei et al., 2023) that focuses on detecting 254 different event types in four categories (industry-related, macrography-related, market-related, and company-related), our collected and organized data further includes the task of identifying causal relationships between events (原因事件类型/Cause event type→结果事件类型/Result event type).
For ZH-PRE, we focus on the Stock Prediction (FinSP) task. It involves categorizing stocks into three categories ("表现不佳(underperforming)", "跑赢大盘(outperforming the market)", "中性(neutral)") by combining stock data and financial news. We collected the StockA dataset (Zou et al., 2022), features stock-specific news and factors for China’s A-shares market, including minute-level price history, setting it apart from similar datasets (Xu and Cohen, 2018; Zhou et al., 2021) by offering comprehensive data on individual stocks.
For ZH-GEN, we include the Text Summarization (FinTS) task. The goal is to distill intricate financial details into succinct summaries. We curated the FinNA dataset (Lu et al., 2023), known as NA, by extracting information from research reports and specifically honing in on their abstracts.
3.2.2 English Financial Datasets
For English, we utilize datasets from PIXIU Xie et al. (2023b), a comprehensive source for English financial data tailored for FinLLMs. The dataset includes five task in financial NLP: classification (EN-CLS), prediction (EN-PRE), extraction (EN-EXT), reasoning (EN-REA) and data-out-of-field (dof) task not used for training LLMs (EN-DOF).
For EN-CLS, our dataset consists of the following three tasks: sentiment analysis (FinSA) , headline classification (FinHC) and credit classification (FinCC). 1) Sentiment Analysis. We used the Financial Phrase Bank (FPB) (Malo et al., 2014) and FiQA-SA (Maia et al., 2018) dataset. FPB comprises a human-annotated phrase bank for financial and economic texts, while FiQA-SA primarily originates from social data, including microblogs, news. 2) Headline Classification. We utilize the Headlines dataset (Sinha and Khandait, 2021), containing news headlines from 2000 to 2019 related to "gold" and annotated with categories like price movements (up, down, stable), historical and future prices, general trends, and asset comparisons. 3) Credit Classification. It involves assessing a customer’s credit status based on a set of attributes and assigning a credit label of "good" or "bad" to the customer. We employed the two datasets, German (Hofmann, 1994) and Australian (Quinlan, ).
For EN-PRE, we approach Stock Prediction as a binary classification task  (Soun et al., 2022), utilizing social media, financial news, and historical stock price data from the past five days to forecast the next-day close price change rate. Classification criteria are: 1) Close price change rate > 0.50% is labeled "Rise"; 2) Close price change rate < -0.50% is labeled "Fall"; 3) Other cases are labeled as "Hold." We utilize three common English datasets: ACL18 (Xu and Cohen, 2018), CIKM18 (Wu et al., 2018), and BigData22 (Soun et al., 2022).
For EN-EXT, we include the Named Entity Recognition task. We use the CoNLL-2003 dataset  (Alvarado et al., 2015) named as NER including sentences from public financial agreements via U.S. SEC filings. It includes four annotated entity types: Location ("LOC"), Organisation ("ORG"), Person ("PER"), and Miscellaneous ("MISC").
For EN-REA, We include the Question Answering task distinct from Chinese, additionally incorporating stock time series and tables to infer answers. We utilized the datasets FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022). FinQA comprises expert-written financial reports of S&P 500 companies, sourced from FinTabNet (Zheng et al., 2021). ConvFinQA expands on FinQA, allowing multi-turn conversations.
Additionally, we introduce EN-OFT including NER, text summarization (FinTS) and hawkish-dovish classification (FinDC) task. 1) Named Entity Recognition. In NER task, we utilize the FINER-ORD  (Alvarado et al., 2015) dataset, which provides labels for entities such as "PER", "LOC", and "ORG". 2) Text Summarization. For this task, we use two datasets: ECTSUM (Mukherjee et al., 2022) and EDTSUM (Zhou et al., 2021). ECTSUM involves extracting key sentences from texts, while EDTSUM focuses on generating fitting news headlines through abstractive summarization. 3) Hawkish-dovish Classification. FinDC aims to classify sentences from monetary policy texts into a "hawkish" or "dovish" stance, unlike FinSA, using the FOMC  (Shah et al., 2023) dataset.
3.2.3 Translation Datasets
To enhance the model’s bilingual and cross-lingual capabilities, we utilized ChatGPT to translate 8 English datasets (ZH-TRA) spanning 4 tasks: FinSA, FinSP, FinHC, and FinQA, into Chinese. We fully translated datasets for text-centric tasks like FinSA and FinHC, while translating only the validation sets for the more intricate FinSP and FinQA tasks. The English FinSA datasets FPB and FiQA-SA were rendered as CFPB and CFiQA-SA in Chinese, featuring sentiment labels: "消极(Negative)", "中性(Neutral)", and "积极(Positive)". The Headlines dataset was transformed into CHeadlines in Chinese, with labels "是(Yes)" and "否(No)". For QA tasks like EnQA and ConvFinQA, we tailored prompts to preserve vital information such as tables, stock names, resulting in new Chinese datasets named CEnQA and CConvFinQA. In FinSP tasks like ACL18, CIKM18, and BigData22, we employed regular expressions to extract text containing task prompts and stock descriptions. Subsequently, by using human-crafted translation prompts, we translated the text while retaining stock details like stock tickers and historical stock prices, yielding translated Chinese datasets named CACL18, CCIKM18, and CBigData22, respectively. To maintain data quality, we removed error-prone or untranslatable data. Our translation method example is available in Appendix Figure 5.
4 ICE-FIND: Bilingual Financial Instruction Data
Base on the different task of raw datasets, we further construct our bIlingual Chinese-English Financial InstructioN Dataset (ICE-FIND), covering 13 tasks and 36 datasets with 604k data samples in total, as detailed in Appendix Table 6. This dataset compilation combined raw data with expertly devised prompts, generating 30-40 unique prompts for each dataset. However, it is important to note that expert-authored prompts may not always accurately reflect the quality of the generated results. We assessed these prompts through human evaluations  [^6] Examples and thier translations are shown in Appendix Figure 7-8 focusing on accuracy, naturalness, and informativeness using a 1-3 scale, to identify the most effective ones. These prompts were tested on various LLMs using the ChatALL   [^7] https://github.com/sunner/ChatALL platform, showing satisfactory performance in most cases.
To handle prompt variations across LLMs, we adopted two strategies. For English tasks, we matched each data with all prompts except for FinQA and ConvFinQA. For Chinese tasks, we randomly selected a prompt for each data due to the larger scale of the raw data. Finally, we converted the raw datasets into instruction-tuning samples by collecting human-designed instructions with prompts   [^8] Examples are presented in Appendix Table 9, input texts, and responses. The instruction data was structured in JSON format (Figure 3). {data_id} represents the ID of each data, {prompt} is the task-specific prompt, {conext} refers to the input data, {Your answer} denotes the expected answer, and {Your label} represents the truly answer.
Figure 3: The format of the instruction dataset.
5 ICE-INTERN: Bilingual Financial Large Language Model
Upon evaluating performance of various backbone LLMs on ICE-FLAR, we explore InternLM-7B to bulid ICE-INTERN-7B with the BiFinID instruction dataset. To assess the impact of different data types on fine-tuning ICE-INTERN, we gradually incorporate the DLC, DLE, DTT, DTE datasets during the fine-tuning process to create variants of ICE-INTERN named ICE-INTERN-dlc-7B, ICE-INTERN-dle-7B, ICE-INTERN-dtt-7B, and ICE-INTERN-full-7B, respectively. We use the same set of fine-tuning hyper-parameters for all models, which are described below. To improve the model’s capability in handling long text, we employ an advanced fine-tuning technique called QLoRA (Hu et al., 2021) to enhance the training process, leading to optimized memory usage and faster training times. Specifically, the training sequences are uniformly cut to a length of 2048 tokens. Then, we use AdamW optimizer with an initial learning rate 5e-5 on weight decay 1e-5 for all parameters. Additionally, the warmup steps are set to 1% of the total training steps. Finally, we fine-tune model total 1 epoch in batch size of 24 on 8 A100 40GB GPUs. Due to the growing volume of fine-tuning data and constraints on our computational resources, we have opted not to proceed with additional fine-tuning of ICE-INTERN using larger parameters.
6 ICE-FLARE: Bilingual Financial Evaluation Benchmark
Based on ICE-FIND, we design our bilingual multi-task evaluation benchmark, ICE-FLARE. The existing benchmarks ([1] FLARE (Xie et al., 2023b), [2] CFBenchmark (Lei et al., 2023), [3] FinCUGE (Chen et al., 2023), [4] FLUE (Shah et al., 2022), [5] Fineval (Zhang et al., 2023b), [6] CGCE (Zhang et al., 2023c)) compared for ICE-FLARE are listed in Table 2. Unfortunately, they are all designed exclusively for either Chinese or English. ICE-FLARE has highlighted its superiority over FLUE in English through FLARE  (Xie et al., 2023b). From Table 2, it is evident that ICE-FLARE introduce bilingual capabilities, significantly surpasses these benchmarks on task diversity and data scalability. Despite FinCUGE having a higher number of financial evaluation tasks, the difference in dataset size compared to ICE-FLARE is nearly 8 times. These public benchmarks have provided valuable references and inspiration for constructing the test set, making ICE-FLARE more comprehensive in terms of task coverage, metric diversity, and sample quantity. However, they still face challenges in evaluating the bilingual capabilities in Chinese-English and severely lack DOF tasks (such as FinDC) to test the robustness of the models. In conclusion, ICE-FLARE, which offers a bilingual evaluation benchmark and a more comprehensive assessment of diversity tasks, is necessary for improving performance and linguistic capabilities of bilingual FinLLMs.
Specific Task Metric Lang Data Test Cover
[1] [2] [3] [4] [5] [6]
[…]
StockB 1,962
CFPB 970
CFiQA-SA 233
en FPB 970
FiQA-SA 235
Semantic Matching
FinSM
[…]
News Classification
FinNC
[…]
Negative Judgment
FinNJ
F1 Accuracy zh NSP 500 $usym2717 \usym$2714 $usym2717 \usym$2717
Answer Selection
FinAS
[…]
CACL18 511
CBigData18 159
CCIKM18 86
en ACL18 3,720
BigData18 1,472
CIKM18 1,143
Relationship Extraction
FinRE
F1 Accuracy zh RE 1,489 $usym2717 \usym$2714 $usym2717 \usym$2717
Headline Classification
FinHC
[…]
Credit Classification
FinCC
[…]
Hawkish-dovish Classification
FinDC
[…]
CEnQA 133
CConFinQA 237
en EnQA 1,147
ConFinQA 1,490
Entity Recognition
FinER
[…]
20CCKS 9,159
21CCKS 1,400
22CCKS 11,829
Text Summarization
FinTS
[…]
Open Source $usym2714 \usym$$\usym2714$ $usym2714 \usym$2717
Task Total 14 5 6 8 4 1 1
Data Total 95k 43k 4k 11k 4k 5k 0.2k
[…]
7 Experiments
7.1 Evaluation Setup
In addition to the FinLLMs we mentioned earlier, such as CFGPT, DISC-FinLLM, FinMA, we also select the representative and outstanding LLMs for evaluating performance. 1) ChatGPT (OpenAI   [^9] Https://www.openai.com/chatgpt): ChatGPT-3.5 and ChatGPT-4, developed by OpenAI, are widely recognized LLMs with versatile functionalities. 2) LLaMa (Touvron et al., 2023): LLaMa and advanced LlaMa2 are powerful LLMs developed by Meta AI, with parameter sizes ranging from 7 billion to 650 billion. 3) Baichuan (Yang et al., 2023a): It is a LLM developed by Baichuan-inc for Chinese and English NLP tasks, available in Baichuan-7B and Baichuan-7B-Chat variants. 4) ChatGLM  (Du et al., 2021): It is an advanced bilingual LLM developed by Tsinghua University and ZhiPu AI, with versions like ChatGLM2-6B and ChatGLM3-6B. 5) BLOOM  (Muennighoff et al., 2022): It developed by the BigScience team, is a versatile LLM with parameter sizes ranging from 560M to 176B, including Bloomz-7b1. 6) InternLM (Team, 2023): InternLM is a language model developed by SenseTime, featuring different parameter configurations such as InternLM-7B. 7) Qwen (Bai et al., 2023): It is a series of advanced LLMs introduced by Alibaba Cloud, including Qwen-7B for conversational capabilities in Chinese.
Figure 4: A radar chart of average metrics for different data types (DLC (A), DLE (B), DTT (C), DTE (D), DOF (E)) on their included specific tasks, 5 data types (F), 10 bilingual NLP tasks (G), along with the best result count of all specific-task metrics on both Chinese (ZH) and English (EN) data (H), in general LLMs, FinLLMs, ICE-INTERN variants. Referencing the FinSM’s average metric, take the average the F1 scores on the BQC and AFQMC datasets.
Following  (Wu et al., 2023; Li et al., 2023b; Xie et al., 2023b), we employed two common used zero-shot and few-shot evaluation methods to assess the model’s adaptability and performance.
7.2 Results
We present evaluation results of different LLMs, FinLLMs and ICE-INTERN variants in Figure 4.
As shown in Figure 4, A demonstrates ICE-INTERN’s outstanding performance across various specific tasks, particularly excelling in tasks involving DLC, DLE, and DTT data types, showcasing strong Chinese language capabilities. However, in D-E, it slightly lags behind ChatGPT-4 in English and DOF tasks. It is particularly evident in F. In G, ICE-INTERN excels in 10 Chinese-English bilingual NLP tasks , except in English reasoning and DOF tasks. H further confirms ICE-INTERN ’s overall superiority in handling tasks in both English and Chinese, particularly excelling in Chinese.
We also draw the conclusions from the detailed results in Appendix Table 10 .
Model Overall Performance. Among the total of 63 specific-task metrics, ICE-INTERN variants achieves optimal performance in approximately 68%. Among them, ICE-INTERN-Full, which fine-tuned on the entire dataset, exhibits the highest percentage of all tasks with the best performance. This fully demonstrates the exceptional superiority of ICE-INTERN and highlights the importance of domain instruction fine-tuning in improving the performance of FinLLMs. InternLM overall outperforms LLaMA2, which fully validates our choice of using InternLM instead of LLaMA as the backbone model used in similar work  (Yang et al., 2023c; Wu et al., 2023; Xie et al., 2023b). We observe that that DLC has little impact for ICE-INTERN-dlc-7B. New introduction of DLE further improves the performance of ICE-INTERN-dle-7B significantly. Conversely, introducing DTT for fine-tuning ICE-INTERN-dtt-7B does not result in an improvement and instead leads to decrease. Finally, incorporating DTE for fine-tuning ICE-INTERN-Full-7B demonstrates a substantial improvement. Our discoveries align with prior research (Chen et al., 2023; Xie et al., 2023b), illustrating the efficacy of the InternLM backbone in DLC. Additionally, diverse DLE data types encompassing comprehension and reasoning tasks prove essential for optimizing performance. Moreover, our results underscore the detrimental impact of poor-quality translation datasets on fine-tuned model efficacy. These insights pave the way for innovative advancements in LLM research within the finance domain.
Bilingual Proficienc Analysis. In the cross-lingual tasks of FinHC, FinSA, and FinSP, the ICE-INTERN-Full model exhibits performance improvements ranging from approximately 1% to 10% when incorporating the corresponding translation data in fine-tuning LLMs for the same task. These findings suggest that the impact of supplementing the corresponding translation data on fine-tuning results is not as pronounced for shorter-text tasks (such as FinSA and FinHC) in DLC. However, when dealing with the long-text (including time-series data) task of FinSP in DLC, incorporating the corresponding translation data can significantly enhance the results. This is because short-text specific tasks involves relatively simple grammar and structure of the financial descriptions, and our ICE-INTERN models can already utilize sufficient information in Chinese for answer reasoning. However, when dealing with longer and more complex texts such as those containing time, historical stock prices, and tweets, our ICE-INTERN models benefit from the inclusion of English data, which enhances its ability to reason with contextual information. These observations highlight the importance of considering the length and complexity of the text when determining the effectiveness of incorporating translation data for fine-tuning LLMs in cross-lingual tasks.
Defect Case Analysis. Despite the strengths exhibited by our ICE-INTERN models on known tasks, they under perform relative to ChatGPT-4 on unseen tasks like FinDC, FINER, and FinST in DOF data types, which have not been fine-tuned. This performance gap suggests that a more diverse set of domain-specific tasks is needed for effective fine-tuning. Specifically, in FinST tasks, we utilize a sophisticated prompt design that involves generating label sequences directly. However, our fine-tuned models consistently struggle to produce outputs in the desired formats. Even models like larger-parameter ChatGPT-3.5 and ChatGPT-4, which demonstrate some capability for token labeling in FinER, face difficulties in generating labels and summaries for ECTSUM and EDTSUM in FinST, especially when confronted with longer contextual information. Our conclusion aligns with PIXIU, suggesting that for unfamiliar tasks within current FinLLMs, it will be imperative to furnish more extensive domain-specific tasks and instructional data in the future to further enhance performance.
7.3 Practical Examples
To demonstrate the practical application of ICE-PIXIU, we have provided examples of model response as shown in Appendix Table 11. It can be concluded that ICE-INTERN provides more accurate responses in tasks such as FinED, FinNC, and FinRE, and offers more comprehensive responses in tasks like FinER and FinQA. Particularly in FinER, ICE-INTERN is able to provide all entity names and their corresponding types. Overall, ICE-INTERN outperforms in accuracy and comprehensiveness in financial question answering.
8 Limitations
While this study has positive aspects, we acknowledge limitations: 1) Parameter Constraints: Limited resources hinder training models with over 7B parameters, impacting performance on large datasets. 2) Backbone LLM Adaptability: Fine-tuned model performance varies based on backbone model adaptability to specific tasks, requiring task-specific selection. 3) Prompt Quality Variability: Even with expert annotation, instruction prompts may perform differently across models, affecting model assessment. 4) Risk of Misuse: Open-source models support research but commercial misuse poses financial risks. Regulation is crucial for responsible usage and risk mitigation.
9 Conclusion
In conclusion, we presented innovative cross-lingual ICE-PIXIU financial framework, which consists of the pioneering cross-lingual bilingual financial LLM, ICE-INTERN, and the the evaluation benchmark, ICE-FLARE. The bilingual capability of ICE-INTERN empowers global finance by breaking language barriers, while ICE-FLARE enables comprehensive cross-lingual assessments for various financial specific tasks. The open access nature of ICE-PIXIU promotes collaboration and research in financial NLP. The framework underscores the significance of different types of instruction data in enhancing performance. Moreover, the framework provides a unified solution for diverse financial applications and ensures reliable cross-lingual consistency through rigorous evaluation. Overall, ICE-PIXIU contributes to the progress of research and development in financial NLP, fostering innovation within the field.
Acknowledgement
We would also like to acknowledge the financial support received from the General Program of Applied Basic Research Programs of Science and Technology Department of Yunnan Province (No.202301AT070184), the Science Research Fund Project of Education Department of Yunnan Province (No.2023J0019), the Key R&D Project of Hubei Province (No.2021BAA029) and the Open Project Program of Yunnan Key Laboratory of Intelligent Systems and Computing (No.ISC22Y08).
Appendix
A Examples of Context and Answer in Financial Task
Table 3-5 presents the examples of context description and annotated answer in various financial task in Chinese, English and translation datatsets.
FE
Context: 国检集团和华测检测哪个比较漂亮?业务都差不多,都是检测,并且国检持有碳交易股份,两个都买了。
Answer: 金融的报道偏向积极评价。
StockB
Context: 中烟集团金融网站,融通租赁注册100亿!揭牌仪式有望。
Answer: 金融的报道偏向中性评价。
CFPB
Context: 维萨拉在2007年第三季度的净利润从2006年同期的680万欧元(980万美元)降至300万欧元(430万美元)。
Answer: 金融的报道偏向消极评价。
CFiQA-SA
Context: \$CERN-在50和200MA上方整合,这里是很好的长期进入点,止损位于10MA以下-目标区域\$70。
Answer: 金融的报道偏向积极评价。
FPB
Context: Hearst will be able to consolidate about 20% of all Russian market for advertising in press after the purchase.
Answer: Sentiment analysis of financial news is positive.
FiQA-SA
Context: \$SLV-4.44% at 18 now AWFUL, down from 42.50
Answer: Sentiment analysis of financial news is negative.
BQC
Context: 1:两个小时还没有等到确认电话怎么办?明天会继续联系嘛? 2:下次借款是否不需要电话确认
Answer: 两个金融表达语义不是相似的。
AFQMC
Context: 1:借呗还款日当天不能再借款吗 2:蚂蚁借呗要一次性还清才能再借吗
Answer: 两个金融表达语义相似的。
NL
Context: [中航泰达:拟购买包钢节能34%股权]中航泰达公告,公司拟以2.09亿元现金认购包钢节能新增注册资本1.28亿元,同时以2.59亿元受让北方稀土持有标的公司的1.59亿元注册资本。本次交易前,北方稀土直接持有包钢节能100%股权。本次交易完成后,公司将直接持有包钢节能34%的股权。
Answer: 报道涉及中国区域。
NL2
Context: 欧股集体高开,德国DAX30指数涨0.44%,英国富时100指数涨1.29%,法国CAC40指数涨0.7%,欧洲斯托克50指数涨0.63%。
Answer: 报道涉及国际大盘
NSP
Context: 钱宝网张小雷涉嫌集资诈骗罪被检方提起公诉中国基金报。实体:钱宝网;钱宝
Answer: 所给实体含有负面消息。
CHeadlines
Context: 在亚洲早盘,金价略有下跌,市场正在关注美国的数据。
Answer: 标题中提到金价会下降的意见。
Headlines
Context: Gold holds near 3-1/2 week low as investors opt for riskier assets.
Answer: Yes, the headline suggests a downward direction for gold.
FinevalF
Context: 国债的偿债率是指()?
Answer: 答案是C:当年还本付息额占当年财政收入的比重。
QA
Context: 11月9日上午,佛山市人民政府与徐工集团工程机械有限公司签约,将在南海区建设总投资20亿元的广东生产基地项目。
Answer: 建设投资方是徐工集团
CEnQA
Context: 下表比较了花旗五年普通股的累积总回报,该股在纽约证券交易所上市,代码为201cc201d。日期,花旗,标普500,标普金融指数;2012-12-31,100.0,100.0,100.0; 2013-12-31,131.8,132.4,135.6;2014-12-31,137.0,150.5,156.2;2015-12-31,131.4,152.6,153.9;2016-12-31,152.3,170.8,188.9;2017-12-31,193.5,208.1,230.9。
Answer: 五年回报百分比0.935
CConFinQA
Context: 以下是比较2007年和2008年净收入变动的分析,金额(以百万计)。<table class=’wikitable’><tr><td>1</td><td></td><td>金额(以百万计)</td> </tr> <tr> <td>2</td ><td>2007年净收入</td> <td>\$991.1</td> </tr> <tr><td>3</td> <td>零售电价</td> <td>-17.1(17.1)</td> </tr> <tr> <td>4</td> <td>购买的电力容量</td> <td>-12.0(12.0)</td> </tr> <tr> <td>5</td> <td>净批发收入</td> <td>-7.4(7.4)</td> </tr> <tr> <td>6</td> <td>其他</td> <td>4.6</td> </tr> <tr> <td>7</td><td>2008年净收入</td> <td>\$959.2</td> </tr> </table>。
Answer: 2007年的净收入是\$991.1
EnQA
Context: The following table provides a comparison of the accumulated total return of citi common stock over a period of five years, which is listed on the nyse under the ticker symbol 201cc201d. date, citi, s&p 500, s&p financials. 31-dec-2011, 100.0, 100.0, 100.0; 31-dec-2012, 150.6, 116.0, 128.8; 31-dec-2013, 198.5, 153.6, 174.7; 31-dec-2014,206.3, 174.6, 201.3; 31-dec-2015, 197.8, 177.0, 198.2; 31-dec-2016, 229.3, 198.2, 243.4.
Answer: The percent of the growth for s&p financials cumulative total return from 2013 to 2014 is 26.6.
ConFinQA
Context: The following is an analysis comparing the changes in net income between 2007 and 2008. <table class=’wikitable’><tr> <td>1</td> <td> </td> <td> amount (in millions)</td> </tr> <tr><td>2</td> <td>2007 net revenue</td> <td>\$991.1</td> </tr> <tr> <td>3</td><td>retail electric price</td> <td>-17.1 ( 17.1 )</td> </tr> <tr> <td>4</td> <td>purchased power capacity</td> <td>-12.0 ( 12.0 )</td> </tr> <tr> <td>5</td> <td>net wholesale revenue</td><td>-7.4 (7.4)</td> </tr> <tr> <td>6</td> <td>other</td> <td>4.6</td></tr> <tr> <td>7</td> <td>2008 net revenue</td> <td>\$959.2</td> </tr> </table>。
Answer: The net revenue in 2008 is \$959.2.
Table 3: Examples with context description and annotated answer for various dataset included in the financial specific task of Sentiment Analysis(FinSA), Sentiment Matching(FinSM), News Classification(FinNC), Negative Judgment(FinNJ), Headline Classification(FinHC), Answer Selection(FinAS) and Question Answering(FinQA).
StockA
Context: 新北洋公告,公司控股子公司荣鑫科技董事会审议通过了《关于拟申请公司股票在全国中小企业股份转让系统终止挂牌的议案》,具体详见荣鑫科技(证券代码:839288)披露在全国中小企业股份转让系统(以下简称“新三板”)的相关公告。荣鑫科技基于其自身经营发展及战略规划的需要,拟申请在新三板终止挂牌。 日期 开盘价 闭盘价 2021-01-04 34.63 34.9763 2021-01-05 34.63 35.5122 2021-01-06 35.1659 34.0116 2021-01-07 33.8879 32.3048 2021-01-08 32.0987 32.3213.
Answer: 该股票表现不佳
CACL18
Context: \$codi近期的走向趋势如下,open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30;2015-09-16, -0.8, 0.2, -1.3, 1.0, 1.0, -0.9, -0.1, -0.1, -0.3, 0.0, 0.5; 2015-09-17,-1.4, 0.9, -1.7, 1.6, 1.6, -2.0, -1.7, -1.5, -1.8, -1.6, -1.2; 2015-09-18, -0.9, 0.2, -1.5, 0.1, 0.1, -1.6, -1.8, -1.5, -1.8, -1.7, -1.3; 2015-09-21, -0.5, 0.3, -1.8, 0.7, 0.7, -1.5, -2.3, -2.1, -2.4, -2.3, -2.1; 2015 -09-22, 1.3, 3.4, -0.4, -3.1,-3.1,1.7,0.8,0.9,0.9,0.7,1.0;2015-09-23,-0.8,0.7,-0.8,0.7,0.7,1.0, 0.1,0.3,0.2,-0.0,0.2; 2015-09-24,-0.1,1.5,-0.7,0.1,0.1,0.7,0.1,0.1,0.2,-0.1,0.1; 2015-09-25,1.5, 1.9,0.0,-0.9, -0.9,1.0, 1.0, 0.8, 1.0, 0.8, 0.9; 2015-09-28, 1.2, 1.2, -0.8, -0.2, -0.2, 0.5, 1.2, 1.0, 1.2, 0.9, 1.0; 2015-09-29, 1.3, 2.1, -0.3, -1.8, -1.8, 2.0, 3.0, 2.6, 2.8, 2.9, 2.8. \$codi评级为买入, 18.6%华尔街分析师根据平均评级/目标得出的上涨。
Answer: \$codi的收盘价在2015-9-30会上涨
CCIKM18
Context: \$chk近期的走向趋势如下,date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2017-10-05,-0.7,0.7,-0.9,0.7,0.7,0.0,0.6,-1.3,-3.8,-5.6, -7.0; 2017-10-06, 1.4,1.7,-1.9, -2.3, -2.3, 2.0, 2.9, 1.4, -0.8,-2.8,-4.5; 2017-10-09,0.2,0.7, -1.4,-0.2, -0.2, 1.4, 2.7, 1.9,0.1,-2.2,-3.9; 2017-10-10,7.9, 8.2, 0.0, -6.7, -6.7, 6.8, 8.9, 9.0, 7.5, 5.0,3.2; 2017-10-11, -0.8, 0.8, -4.8, 1.0, 1.0, 4.1, 6.6, 7.4, 6.4, 4.2, 2.4; 2017-10-12,1.3,2.1,-1.8,-2.5,-2.5,4.4,8.1,9.7,8.8,6.8,5.3; 2017-10-13, 0.5,1.3,-0.8, 0.8, 0.8,2.0,6.2,8.2,7.8,6.3,4.6。
Answer: \$chk的收盘价在2027-10-19会上涨
CBigData22
Context: \$msft近期的走向趋势如下,date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2020-10-09, -2.1, 0.0, -2.1, 2.5, 2.5, -2.5, -2.8, -3.6, -4.0, -4.0,-2.5 ; 2020-10-12, -1.2, 1.1, -2.1, 2.6, 2.6, -3.9, -4.7, -5.5, -6.1, -6.3, -5.1; 2020-10-13,-0.1,1.1,-1.1,0.7,0.7,-3.0, -4.6, -5.6, -6.4, -6.5, -5.8; 2020-10-14, 1.0, 1.5, -0.8, -0.9, -0.9, -1.2, -3.3,- 4.2, -5.2, -5.5, -5.0; 2020-10-15,-1.2, 0.3, -1.7, -0.5, -0.5,0.2,-2.4,-3.2,-4.3,-4.7,-4.7; 2020-10-16, 0.2, 1.2, -0.2, 0.0, 0.0, 0.6, -1.8, -2.8, -3.8, -4.5, -4.6; 2020-10-19, 2.9, 3.8, -0.2, -2.5, -2.5, 2.4, 0.9, -0.2, -1.1, -1.9, -2.2; 2020-10-20, 0.5, 1.3, -0.7, 0.2, 0.2, 1.5, 1.1, -0.2, -1.2, -2.0, -2.2; 2020-10-21,-0.8, 1.0, -0.8, 0.1, 0.1, 0.8, 1.2, -0.1, -0.9, -1.8, -2.2; 2020-10-22, -0.4, 0.5, -1.5, 0.0, 0.0, 0.4, 1.4, -0.1, -0.7, -1.7, -2.1。
Answer: \$msft的收盘价在2020-10-23会上涨
ACL18
Context: The recent trend of \$codi is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2015-03-30, 0.0, 0.4, -0.9, 0.0, 0.0, -0.0, -0.8, -1.6, -2.1, -2.4, -2.6, 2015-03-31, 0.1, 0.3, -0.8, -0.3, -0.3, 0.2, -0.4, -1.0, -1.7, -2.0, -2.2; 2015-04-01, 1.1, 1.1, -0.7, -1.2, -1.2, 1.2, 0.9, 0.5, -0.4, -0.7, -0.9; 2015-04-02, -0.2, 0.4, -0.4, -0.2, -0.2, 1.0, 1.2, 0.7, -0.1, -0.4, -0.6; 2015-04-06, -0.6, 0.5, -0.9, 0.6, 0.6, 0.2, 0.6, 0.2, -0.6, -0.8, -1.1;2015-04-07, -0.4, 1.4, -0.4, 0.2, 0.2, -0.2, 0.4, 0.0, -0.5, -1.0, -1.2; 2015-04-08, 0.0, 0.2, -0.4, 0.1, 0.1, -0.4, 0.2, 0.0, -0.4,- 1.0, -1.2; 2015-04-09, -0.5, 0.2, -1.1, 0.6, 0.6, -0.7, -0.4, -0.5, -0.7, -1.4, -1.7; 2015-04-10, 0.1, 0.4, -0.5, 0.2, 0.2, -0.5, -0.6, -0.5,-0.8, -1.5, -1.7;2015-04-13, 1.4, 1.9, 0.0, -0.7, -0.7, 0.2, 0.1, 0.3, -0.0, -0.6, -0.9. \$codi - current report filing (8-k) 2015-04-13: william blair starts compass diversified \$codi at outperform.
Answer: The closing price of \$codi will rise at 2015-04-14.
CIKM18
Context: The recent trend of \$aal is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2017-01-17, 1.7, 2.1, -0.2, -1.9, -1.9, 2.5, 0.9, 1.3, 2.0, 2.3, 2.0; 2017-01-18, -0.5, 0.3, -2.1, 1.9, 1.9, 0.2, -0.7, -0.7, 0.1, 0.2, 0.2; 2017-01-19, 0.8, 1.8, -0.8, -0.8, -0.8, 0.5, 0.2, -0.1, 0.8, 1.0, 1.1; 2017-01-20, -1.0, 0.3, -1.6, 1.6, 1.6, -1.1, -0.9, -1.6, -0.9, -0.6, -0.3; 2017-01-23, 2.0, 2.4, -0.4, -2.2, -2.2,0.8,1.5,0.6,1.1,1.6,1.8 2017-01-24, -1.0, 0.5, -1.6, 1.3, 1.3, -0.1, 0.3, -0.5, -0.3, 0.2, 0.5; 2017-01-25,-0.0, 0.6, -0.6, 0.8, 0.8, -0.8, -0.6, -1.1, -1.1, -0.5, -0.4; 2017-01-26, -2.3, 0.2, -2.5, 3.5, 3.5, -3.2, -3.7, -4.1, -4.4, -3.8, -3.7 2017-01-27,6.5,6.5,-0.5,-5.3,-5.3, 1.8, 1.4, 1.5, 0.9, 1.4, 1.7; 2017-01-30, 1.6, 2.3, -2.3, -4.4, -4.4, 5.5, 5.5, 5.9, 5.3, 5.6, 6.1.
Answer: The closing price of \$aal will fall at 2020-9-28.
BigData22
Context: The recent trend of \$intc is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2020-09-14, -1.1, 0.6, -1.1, 0.3, 0.3, -0.4, 1.3, 1.0, 0.5, 0.1, -0.2; 2020-09-15, -0.4, 1.2, -0.5, 1.2, 1.2, -1.1, -0.1, -0.1, -0.6, -1.0, -1.3; 2020-09-16, 0.3, 1.3, -0.4, 0.7, 0.7, -1.5, -0.9, -0.7, -1.2, -1.5, -1.9; 2020-09-17, -1.9, 0.3, -2.0, -0.1, -0.1, -0.9, -1.2, -0.5, -0.9, -1.3, -1.7; 2020-09-18, 0.9, 1.2, -1.7, -0.9, -0.9, 0.2, -0.4, 0.4, 0.1,-0.4,-0.7; 2020-09-21, -0.7, 0.0, -1.8, -0.3, -0.3, 0.7, -0.1, 0.7, 0.5, 0.0, -0.3; 2020-09-22, -0.1, 0.5, -1.0, 0.5, 0.5, 0.2, -0.4, 0.1, 0.1, -0.4, -0.7; 2020-09-23, 2.1, 2.7, -0.3, -2.3, -2.3, 1.9, 1.7, 2.1, 2.3, 2.0, 1.6;2020-09-24, -1.3, 1.0, -1.5, 0.7, 0.7, 0.7, 1.1, 1.0, 1.6, 1.3, 0.9; 2020-09-25, -2.0, 0.7, -2.4, 1.6, 1.6, -0.8, -0.4, -0.6, 0.0, -0.2, -0.5.
Answer: The closing price of \$intc will rise at 2020-9-28.
Table 4: Examples with context description and annotated answer for various dataset included in the financial specific task of Stock Prediction(FinSP).
19CCKS
Context: 恒立实业(000622)遭证监会立案调查 涉嫌信披违规上海家化(600315)复星退出家化集团股权竞购。
Answer: 事件类型是信披违规,事件主体为恒立实业
20CCKS
Context: 2018年10月,中弘股份于发布公告称,截至2018年9月28日,公司逾期债务本息合计金额超过55亿元,全部为各类借款。
Answer: 事件类型是债务违约,事件主体是中弘股份
21CCKS
Context: 反观需求,由于生猪存栏量持续下降,而水产养殖启动缓慢,豆粕需求难以有效放大。供需失衡,导致国内外豆粕价格在豆类板块中呈现弱势。
Answer: 原因事件类型是供给减少,原因实体是生猪, 结果事件类型是市场价格下降,结果实体是豆粕
22CCKS
Context: 易投配资为什么无法出金;晨曦航空股东高文舍拟减持不超3%股份;招行逾期第三个月的情况。
Answer: 事件类型是股东减持,事件主体是晨曦航空
RE
Context: 张裕与卡斯特集团合作的牵线人陈光因嫌分红太少而有意退出。实体对:张裕-卡斯特集团
Answer: 张裕和卡斯特集团的关系为合作
NA
Context: 美今日凌晨,苹果发布iOS15.4首个测试版,更新后 Face ID将支持戴口罩解锁。 按照官方描述,在升级到戴口罩使用面容ID之后,开机就会提醒是否要使用此功能,用户选择使用后,会要求再录入一次面容ID,这次主要是记录使用者的眼睛周围特征,用于戴口罩时候识别。
Answer: 文本的摘要为苹果iOS15.4测试版支持戴口罩的面容解锁。
NER
Context: The proceeds of the equipment advances will be used solely to reimburse Borrower for the purchase of eligible equipment.
Answer: Entity name is Borrower, and the corresponding entity type is PER.
CNER
Context: 数、数年数月数日,银保监会核准王小林先生本行董事任职资格。
Answer: 实体名称是银保监会, 类型是ORG; 实体名称是王小林, 类型是PER
FINER-ORD
Context: McPhail , 50 , is angry and perplexed by politicians clamouring to sign a free trade deal with China that allows the importation of Chinese workers .
Answer: The "token:label" labeling result for each row of data is [Bill:B-PER Shorten:I-PER pictured:O with:O the:O ALP:B-ORG ’s:O candidate:O for:O Canning:B-LOC ,:O Matt:B-PER Keogh:I-PER .:O]
ECTSUM
Context: And I’m confident that our disciplined approach to operating the business will result in our continued success throughout the balance of fiscal 2021. Coupled with the cost reduction actions we recently implemented, …, Our Test business delivered a really solid Q1 by significantly beating our internal expectations and delivering an EBITDA margin of nearly 13% versus 11% last Q1.
Answer: The label whether a sentence is included in the abstract is [1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0].
EDTSUM
Context: ORLANDO, Fla., June 4, 2020 PRNewswire–Darden Restaurants, Inc., (NYSE: DRI) plans to release its fiscal 2020 fourth quarter financial results…For more information, please visit www.darden.com. Logo: https:mma.prnewswire.com media 340562 darden-restaurants-inc-logo.jpg SOURCE Darden Restaurants, Inc.: Financial Related Links http:www.Darden.com
Answer: The generated summary is Darden Restaurants to Host Fiscal 2020 Fourth Quarter Conference Call on June 25.
German
Context: The client has attributes: X1: A14, X2: 12, X3: A32, X4: A40, X5: 1386, X6: A63, X7: A73, X8: 2, X9: A92, X10: A101, X11: 2, X12: A122, X13: 26, X14: A143, X15: A152, X16: 1, X17: A173, X18: 1, X19: A191, X20: A201.
Answer: The user’s creditworthiness is bad.
Australian
Context: The client has attributes: A1: 0.0, A2: 68.67, A3: 15.0, A4: 2.0, A5: 10.0, A6: 9.0, A7: 0.0, A8: 1.0, A9: 1.0, A10: 14.0, A11: 0.0, A12: 2.0, A13: 0.0, A14: 3377.0.
Answer: good is the user’s creditworthiness.
FOMC
Context: Nonetheless, given a recovery in U. S. domestic demand approximating their current forecasts, growth in imports likely would exceed that of exports by a wide margin over the forecast horizon.
Answer: The stance of this monetary policy text is hawkish.
Table 5: Examples with context description and annotated answer for various dataset included in the financial specific task of Event Detection(FinED), Relationship Extraction(FinRE), Text Summarization(FinTS), Entity Recognition(FinER), Credit Classification(FinCC) and Hawkish-dovish Classification(FinDC).
B Details Information of Raw Data Statistics
Table 6 presents the specifics of the raw data used to construct the Chinese-English bilingual dataset for multi-task financial instruction tuning.
Lang D_T NLP_T S_T Dataset Raw Instruction Evaluation Data Source License
ZH DLC ZH-CLS FinSA FE 18,177 18,177 2,020 social texts Public
StockB 9,812 9,812 1,962 social texts Apache-2.0
FinSM BQC 120,000 110,000 10,000 bank service logs Public
AFQMC 38,650 38,650 4,316 online chat service Apache-2.0
FinNC NL 7,955 7,955 884 news articles Public
NL2 7,955 7,955 884 news articles Public
FinNJ NSP 4,499 4,499 500 social texts Public
FinAS FinevalF 1,115 1,115 222 financial exam Apache-2.0
FinRE RE 14,973 14,973 1,489 news, entity pairs Public
ZH-PRE FinSP StockA 14,769 14,769 1,477 news,historical prices Public
DLE ZH-EXT FinQA QA 22,375 22,375 2,469 QA pairs of news Public
FinER CNER 1,685 1,685 337 financial reports Public
FinED 19CCKS 156,834 14,674 2,936 social texts CC BY-SA 4.0
20CCKS 372,810 45,796 9,159 news, reports CC BY-SA 4.0
21CCKS 8,000 7,000 1,400 news, reports CC BY-SA 4.0
22CCKS 109,555 59,143 11,829 news, reports CC BY-SA 4.0
ZH-GEN FinTS NA 32,400 32,400 3,600 news, announcements Public
DTT ZH-TRA FinSA CFPB 4,845 4,838 970 economic news MIT license
CFiQA-SA 1,173 1,143 233 ews headlines,tweets MIT license
FinSP CACL18 27,056 2,555 511 tweets, historical prices MIT license
CBigData22 7,167 798 159 tweets, historical prices MIT license
CCIKM18 4,970 431 86 tweets, historical prices MIT license
FinHC CHeadlines 102,708 10,256 2,051 news headlines MIT license
FinQA CEnQA 8,281 668 133 earnings reports MIT license
CConFinQA 12,594 1,189 237 earnings reports MIT license
EN DTE EN-CLS FinSA FPB 4,845 4,845 970 economic news CC BY-SA 3.0
FiQA-SA 1,173 1,173 235 news headlines,tweets Public
FinHC Headlines 11,412 102,708 20,547 news headlines CC BY-SA 3.0
FinCC German 1,000 1,000 200 credit records CC BY-SA 4.0
Australian 690 690 139 credit records CC BY-SA 4.0
EN-PRE FinSP ACL18 27,053 27,053 3,720 tweets, historical prices MIT License
BigData22 7,164 7,164 1,472 tweets, historical prices Public
CIKM18 4,967 4,967 1,143 tweets, historical prices Public
EN-EXT FinER NER 609 609 98 financial agreements CC BY-SA 3.0
EN-REA FinQA EnQA 8,281 8,281 1,147 earnings reports MIT License
ConFinQA 3,458 12,594 1,490 earnings reports MIT License
DOF EN-DOF FinER FINER-ORD 1,075 - 1,075 news articles CC BY-SA 4.0
FinTS ECTSUM 495 - 495 earning call transcipts Public
EDTSUM 2,000 - 2,000 news articles Public
FinDC FOMC 496 - 496 FOMC transcripts CC BY-SA 4.0
Table 6: The detailed information regarding the raw data for the Chinese-English bilingual multi-task financial instruction and evaluation encompasses language proficiency (Lang), data type (D_T), NLP task (NLP_T), specific task (S_T), dataset name, instruction data size, evaluation data size, data source, and license information.
C Example of Chinese Dataset Translation
Figure 5 presents the example of translating BigData22 into Chinese CBigData22 Chinese dataset using human-written specific translation prompts.
Figure 5: Giving the example of translating BigData22 English dataset into CBigData22 Chinese dataset with human-written specific translation prompts. Specifically, we will translate the task prompts and twitter descriptions from BigData22 separately, while preserving the original stock information, which includes stock tickers with \$ signs and historical stock prices, unchanged.
D Human Evaluation for Human-written Prompts
Table 7-8 presents the example of human evaluation for domain-expert written prompts on various financial specific task
Specific Prompt Input Target Response Metric
Task acc nat inf AVG
Sentiment Analysis (FinSA)
确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。
涨空间异常巨大,中铁装配,实在是未来的大牛股啊
积极
积极
3 3 3 3
Sentiment Matching (FinSM)
你需要判断给出的两个金融文本表达的意思是否相似,请回答是或否。
花呗咋扫不上/花呗扫一扫不能付钱
0 1 1 0.7
News Classification (FinNC)
请根据金融报道的内容,判断其涉及范围属于’中国’、’国际’还是’外国’?
伯恩斯坦:首予康菲石油(COP.N)跑赢大盘评级,目标价为98美元。
外国
国际
1 3 1 1.7
Negative Judgment (FinNJ)
根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’。
自从号称750亿交易量的唐小僧倒下…几乎每天都有几家网贷公司宣布清盘、提现困难、经侦介入等。实体:唐小僧;多米
3 3 3 3
Answer Selection (FinAS)
根据所提出的金融问题,请从以下四个选项中选择最合适的一个。你的输出应该是:’A’、’B’、’C’ 或’D’。
如果一个竞争性市场位于长期均衡状态中,那么所有的厂商()。 A,采用完全相同的生产工艺; B,具有同一最低平均的成本; C,都能获得经济利润; D,以上全对
D
D
3 3 2 2.7
Stock Prediction (FinSP)
请根据新闻对股票数据的影响判断该公司的股票运动走势是跑赢大盘、中性还是表现不佳。
高新兴公告,实控人拟减持不超过6%。日期 开盘价 闭盘价,2020-09-23, 5.56, 5.52;2020-09-24, 5.47, 5.22; … ; 2020-09-29, 5.02, 5.03
表现不佳
无法预测
0 1 2 1
Headline Classification (FinHC)
下面所给内容是否提到了黄金商品的上涨趋势?请回答是或否。
金价在印度的主要城市上涨
3 3 2 2.7
Question Answering (FinQA)
你需要分析金融文本,根据内容回答相关问题。
【亿邦动力讯】10月21日消息,洛丁森完成B轮融资,…,下文中发生的事件类型有哪些?
企业融资
无法确定
0 0 0 0
Event Detection (FinED)
阅读以下金融领域的公告,判断所有的事件类型及其对应主体。请以:’事件类型, 事件主体’的格式回复。其中事件类型应该在[’信批违规’,…,’涉嫌欺诈’]中。
劲嘉股份(002191)因虚假陈述遭证监会处罚浩云科技(300448)股东披露减持计划
信批违规, 劲嘉股份
信批违规, 劲嘉股份
3 3 3 3
Entity Recognition (FinER)
在下面所给金融语句中,识别指明个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体。
报告期内,本公司除中国银保监会批准的经营范围内的金融担保业务外,没有其他需要披露的重大担保事项。
中国银保监会, ORG
中国银保监会, ORG;中国, LOC
2 2 3 2.3
Relationship Extraction (FinRE)
请仔细分析所给金融报道和实体对,然后在[’合并’, ’竞争’,…]中选择能准确描述该实体对关系的选项。
张裕与卡斯特集团合作的牵线人陈光因嫌分红太少而有意退出。实体对:卡斯特集团-张裕
合作
合作关系受损
1 2 3 2
Text Summarization (FinTS)
你的任务是对所给的金融文本进行一句话摘要,要求尽量简短。
前段时间,佳能宣布关闭经营了32年的珠海工厂,引发网络关注。…而优厚的补偿,也是佳能自愿选择的“体面离开”。(新京报)
佳能撤离珠海,不宜说成“恶意赔偿”
佳能关闭珠海工厂,提供优厚补偿,引发争议。
1 2 3 2
Table 7: Giving the example of human evaluation (including accuracy(acc), naturalness(nat), informativeness(inf), and average(AVG) metrics) for domain-expert written prompts on various financial specific task. We score (1-3) for all metrics and retain the high-quality (can effectively provide target format on ChatAll) prompts with AVG above 2.
Specific Prompt Input Target Response Metric
Task acc nat inf AVG
Sentiment Analysis (FinSA)
Determine the mood of the sentences in the financial news article provided, identifying whether the mood in them is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’.
Up space is exceptionally huge, China Railway assembly, is really the future of the big bull stock ah
positive
positive
3 3 3 3
Sentiment Matching (FinSM)
You need to decide whether the two financial texts given express similar meanings, answer yes or no.
Why can’t I pay/ I can’t pay by scanning.
yes
no
0 1 1 0.7
News Classification (FinNC)
Based on the content of the financial report, please determine whether its scope is ’Chinese’, ’international’ or ’foreign’?
Bernstein: first Outperform rating on ConocoPhillips (COP.N), \$98 price target.
foreign
international
1 3 1 1.7
Negative Judgment (FinNJ)
Based on the given financial news and entity, you need to determine whether the given entity contains negative news or not, you simply need to answer ’yes’ or ’no’.
Since the fall of Tang Xiaosheng, which claimed a transaction volume of 75 billion … Almost every day, there are several online lending companies announced liquidation, cash withdrawal difficulties, the investigation intervention, etc. Entity: Tang Xiaosheng; Domi
yes
yes
3 3 3 3
Answer Selection (FinAS)
Based on the financial question posed, choose the most appropriate of the following four options. Your output should be: ’A’, ’B’, ’C’ or ’D’.
If a competitive market is in long-run equilibrium, then all firms () A, use exactly the same production process; B, have the same minimum average cost; C, are economically profitable; D, all of the above.
D
D
3 3 2 2.7
Stock Prediction (FinSP)
Determine whether the company’s stock movement trend is outperforming, neutral, or underperforming based on the impact of the news on the stock data.
Gaoxin announced that the real controller intends to reduce its holdings by no more than 6%. Date Opening price Closing price : 2020-09-23, 5.56, 5.52 ; 2020-09-24, 5.47, 5.22; … ; 2020-09-29, 5.02, 5.03
underperformance
unpredictable
0 1 2 1
Headline Classification (FinHC)
Does the content given below refer to an upward trend in gold commodities? Please answer yes or no.
Gold prices rise in India’s major cities
yes
yes
3 3 2 2.7
Question Answering (FinQA)
You will need to analyze financial texts and answer relevant questions based on the content.
[YPC] October 21, 2011 - Lodgingsen closed its Series B financing… What are the types of events that occur in the following article?
corporate financing
indeterminate
0 0 0 0
Event Detection (FinED)
Read the following announcements in the financial sector and determine all event types and their corresponding subjects. Please respond in the format :’Event Type, Event Subject’. Where the event type should be in [’credit approval violation’, …’ Suspected Fraud’].
Jinjia shares (002191) due to misrepresentation by the Securities and Futures Commission penalized haoyun science and technology (300448) shareholders disclosure of shareholding reduction plan
Letter approval violation, Jinjia shares
Letter approval violation, Jinjia shares
3 3 3 3
Entity Recognition (FinER)
In the financial statement given below, identify the specific named entity that specifies a person (’PER’), organization (’ORG’), or location (’LOC’).。
During the reporting period, the Company did not have any other material guarantee matters requiring disclosure, except for the financial guarantee business within the scope of operation approved by the China Banking and Insurance Regulatory Commission.
China Banking and Insurance Regulatory Commission, ORG
China Banking Regulatory Commission, ORG; China, LOC
2 2 3 2.3
Relationship Extraction (FinRE)
Carefully analyze the given financial reports and entity pairs, and then among [’Merger’, ’Competition’, …] Choose the option that accurately describes the relationship of the entity pair.
Chen Guang, the matchmaker of Zhangyu’s partnership with Custer Group, intends to quit because he thinks the dividend is too small. Entity Pair: Custer Group - Zhang Yu
collaborate
Impaired partnership
1 2 3 2
Text Summarization (FinTS)
Your task is to make a one-sentence summary of the given financial text, keeping it as short as possible.
Some time ago, Canon announced the closure of its Zhuhai plant, which had been in operation for 32 years, triggering concern on the Internet…. The generous compensation is also Canon’s voluntary choice to "leave with dignity". (New Beijing News)
Canon’s withdrawal from Zhuhai should not be labeled as "malicious compensation".
Canon closes Zhuhai plant, offers generous compensation, sparks controversy.
1 2 3 2
Table 8: Giving the example of human evaluation (including accuracy(acc), naturalness(nat), informativeness(inf), and average(AVG) metrics) for domain-expert written prompts on various financial specific task. We score (1-3) for all metrics and retain the high-quality (can effectively provide target format on ChatAll) prompts with AVG above 2.
E Examples of Instruction Prompt in Chinese Task
Table 9 presents the example of the Chinese prompt and corresponding English translation for each specific financial task.
Specific Task Chinese Prompt English Translation
Sentiment Analysis (FinSA)
确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。
Determine the sentiment of the sentences in the provided financial news article, identifying whether the sentiment is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’.
Sentiment Matching (FinSM)
判断给出的两个金融文本表达的意思是否相似,你只需要回答是或否。现在判断下面两个句子是否相似:
Determine whether the two provided financial texts convey similar meanings; you only need to answer ’yes’ or ’no’. Now determine if the following two sentences are similar.
News Classification (FinNC)
根据金融报道的内容,判断其属于’中国’、’国际’还是’外国’,只需给出其中一个类别。现在分析下面的报道,回答它所属的类别:
According to the content of the financial report, determine whether it belongs to ’China’, ’International’, or ’Foreign’, only one category needs to be given. Now analyze the following report and answer the category it belongs to:
Negative Judgment (FinNJ)
根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’
Based on the given financial news and entities, you need to determine whether the given entities contain negative messages. You only need to simply answer ’yes’ or ’no’.
Answer Selection (FinAS)
根据所提出的金融问题,从以下四个选项中选择最合适的一个。你的输出应该是:’A’、’B’、’C’ 或’D’。
Based on the financial issue presented, choose the most appropriate one from the following four options. Your output should be: ’A’, ’B’, ’C’, or ’D’.
Stock Prediction (FinSP)
请仔细分析数据和推文,预测2017-10-13时\$trv的收盘价格是上涨还是下跌。请确认是上涨还是下跌, 只回答为’上涨’或者’下跌’。
Please carefully analyze the data and tweets, and predict whether the closing price of \$TRV on 2017-10-13 will go up or down. Please confirm whether it is going up or down. You can answer ’Rise’ or ’Fall’.
Headline Classification (FinHC)
请考虑标题是否讨论了与黄金相关的过去事件。新闻标题是否暗示黄金的过去新闻?您的回答应该为’是’或’否’。
Please consider whether the headline discusses past events related to gold. Does the news headline imply past news about gold? Your answer should be ’Yes’ or ’No’.
Question Answering (FinQA)
你需要分析金融文本,根据内容回答相关问题。如果你对答案感到迷茫,可以回答’无相应参数’。
You need to analyze financial texts and answer related questions based on the content. If you feel confused about the answer, you can reply with ’No relevant parameters’.
Event Detection (FinED)
阅读以下金融领域的公告,判断所有的事件类型及其对应主体,省略数字。请以:’事件类型, 事件主体’的格式回复 。其中事件类型应该在这里面:[’信批违规’,…,’涉嫌欺诈’]
The task is to read a financial announcement, determine all event types and their corresponding entities, and reply in the format ’event type, event subject’. The event types should be among those listed in the brackets [‘credit approval violation’,…,’suspected fraud’].
Entity Recognition (FinER)
在分析中国证券交易委员会备案文件中的句子时,识别指明个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体。答案应遵循格式’实体名称, 实体类型’ 。
When analyzing sentences in the China Securities Regulatory Commission filing documents, identify specific named entities that indicate individuals (’PER’), organizations (’ORG’), or locations (’LOC’). The answer should follow the format ’entity name, entity type’.
Relationship Extraction (FinRE)
请仔细分析所给金融报道和实体对,然后在[’合并’, ’竞争’,…]中选择能准确描述该实体对关系的选项。请直接给出答案,如有疑虑可回答unknown。
Please carefully analyze the given financial report and entity pair, then choose the option from [’merge’, ’compete’, …] that accurately describes the relationship of the entity pair. Please provide the answer directly; if in doubt, you may respond with ’unknown’
Text Summarization (FinTS)
通过阅读金融公告,你的任务是对所给的文本进行简短的总结,重点突出主要论点,长度保持在一到两句之间。
Your task is to provide a brief summary of the given text by reading financial announcements, emphasizing the main arguments, and keeping the length between one to two sentences
Table 9: Examples of the Chinese prompt and corresponding English translation for various specific financial task both in Chinese and translation datasets. Each specific task has expert 10-20 diverse prompts annotated and evaluated by domain experts, and we randomly select one as an example for demonstration.
F Evaluation Results of LLMs, FinLLMs and ICE-INTERN variants
Table 10 presents the detailed evaluation results of different 6B-7B LLMs, baseline FinLLMs and variants of ICE-INTERN on the ICE-FLARE benchmark.
Language
Data Type
NLP Task
Specific Task
Metric
Dataset
Bloomz-7B1
Baichuan-7B
Baichuan2-7B
LLaMa2-6B
ChatGLM2-6B
ChatGLM3-6B
Qwen-7B
ChatGPT
GPT-4
InternLM-7B
CFGPT-sft-7B
DISC-FinLLM-13B
FinMA
ICE-INTERN-dlc-7B
ICE-INTERN-dle-7B
ICE-INTERN-dtt-7B
ICE-INTERN-full-7B
ZH DLC ZH-CLS FinSA Accuracy FE 0.587 0.442 0.002 0.423 0.000 0.572 0.366 0.329 0.349 0.553 0.306 0.624 0.534 0.760 0.778 0.743 0.760
StockB 0.196 0.058 0.000 0.047 0.000 0.389 0.431 0.310 0.441 0.297 0.070 0.282 0.202 0.422 0.593 0.636 0.593
F1 StockB 0.177 0.098 0.000 0.078 0.000 0.443 0.482 0.313 0.462 0.279 0.098 0.344 0.201 0.395 0.571 0.575 0.578
FinSM F1 BQC 0.008 0.154 0.002 0.292 0.002 0.444 0.334 0.682 0.704 0.278 0.218 0.255 0.230 0.854 0.859 0.846 0.854
AFQMC 0.004 0.296 0.014 0.266 0.012 0.470 0.560 0.323 0.391 0.292 0.460 0.373 0.095 0.743 0.609 0.594 0.743
FinNC F1 NL 0.245 0.219 0.002 0.280 0.000 0.453 0.348 0.645 0.816 0.375 0.164 0.267 0.261 0.950 0.922 0.954 0.950
NL2 0.000 0.049 0.000 0.066 0.000 0.049 0.074 0.034 0.273 0.039 0.004 0.002 0.060 0.786 0.774 0.768 0.786
FinNJ F1 NSP 0.041 0.217 0.000 0.353 0.004 0.558 0.636 0.935 0.916 0.351 0.435 0.208 0.442 0.692 0.692 0.695 0.692
FinAS Accuracy FinevalF 0.302 0.329 0.014 0.266 0.185 0.423 0.243 0.491 0.662 0.320 0.284 0.369 0.261 0.441 0.410 0.387 0.441
F1 FinevalF 0.295 0.319 0.024 0.171 0.190 0.418 0.096 0.481 0.686 0.247 0.264 0.362 0.220 0.442 0.387 0.376 0.442
FinRE F1 RE 0.012 0.039 0.000 0.011 0.000 0.076 0.026 0.211 0.340 0.015 0.014 0.108 0.009 0.092 0.088 0.103 0.103
ZH-PRE FinSP F1 StockA 0.131 0.000 0.000 0.000 0.000 0.220 0.154 0.422 0.383 0.003 0.050 0.315 0.000 0.340 0.637 0.630 0.617
DLE ZH-CLS FinQA Accuracy QA 0.098 0.001 0.000 0.001 0.000 0.099 0.000 0.256 0.064 0.001 0.092 0.201 0.123 0.167 0.864 0.860 0.855
FinER Entity F1 CNER 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.101 0.103 0.000 0.000 0.000 0.000 0.000 0.483 0.482 0.468
FinED Precision 19CCKS 0.080 0.013 0.003 0.001 0.012 0.001 0.001 0.010 0.011 0.003 0.003 0.004 0.004 0.005 0.831 0.825 0.824
20CCKS 0.052 0.006 0.002 0.001 0.009 - 0.001 0.001 0.008 0.002 0.002 0.002 0.003 0.006 0.740 0.741 0.743
21CCKS 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.001 0.004 0.000 0.000 0.000 0.000 0.000 0.001 0.001 0.002
22CCKS 0.000 0.001 0.001 0.000 0.002 0.001 0.000 0.006 - 0.001 0.003 0.000 0.021 0.007 0.647 0.642 0.660
F1 19CCKS 0.081 0.013 0.003 0.001 0.012 0.001 0.001 0.010 0.012 0.003 0.006 0.005 0.007 0.005 0.831 0.825 0.824
20CCKS 0.068 0.011 0.002 0.003 0.015 - 0.002 0.001 0.010 0.003 0.004 0.004 0.005 0.004 0.705 0.703 0.702
21CCKS 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.001 0.003 0.000 0.000 0.000 0.000 0.000 0.001 0.001 0.001
22CCKS 0.001 0.001 0.001 0.001 0.004 0.001 0.001 0.008 - 0.001 0.004 0.000 0.002 0.007 0.588 0.581 0.595
ZH-GEN FinTS Rouge-1 NA 0.078 0.028 0.006 0.104 0.001 0.256 0.142 0.093 - 0.145 0.006 0.014 0.010 0.050 0.284 0.289 0.287
BertScore NA 0.676 0.527 0.482 0.649 0.496 0.748 0.603 0.710 - 0.651 0.603 0.607 0.603 0.707 0.824 0.823 0.824
BartScore NA -5.704 -6.845 -7.274 -4.539 -7.293 -4.097 -4.301 -5.882 - -4.153 -6.661 -6.194 -6.640 -6.044 -4.180 -4.183 -4.027
DTT ZH-TRA FinSA Accuracy CFPB 0.005 0.203 0.000 0.098 0.000 0.424 0.239 0.699 0.196 0.273 0.256 0.085 0.038 0.545 0.498 0.846 0.849
F1 CFPB 0.010 0.129 0.000 0.092 0.000 0.348 0.340 0.642 0.306 0.209 0.346 0.091 0.067 0.529 0.485 0.845 0.846
CFiQA-SA 0.652 0.401 0.000 0.350 0.000 0.617 0.773 0.400 0.120 0.668 0.420 0.784 0.616 0.752 0.709 0.796 0.868
FinSP Accuracy CACL18 0.550 0.585 0.591 0.577 0.589 0.485 0.556 0.544 0.568 0.566 0.581 0.564 0.585 0.438 0.442 0.581 0.417
CBigData18 0.667 0.686 0.686 0.686 0.686 0.440 0.673 0.629 0.642 0.686 0.535 0.585 0.660 0.365 0.403 0.660 0.321
CCIKM18 0.384 0.372 0.372 0.372 0.372 0.605 0.384 0.326 0.419 0.372 0.407 0.477 0.361 0.605 0.628 0.395 0.605
MCC CACL18 -0.034 0.008 0.000 0.021 -0.002 0.004 0.012 0.002 0.010 -0.017 0.099 0.076 0.022 0.044 0.042 0.039 -0.034
CBigData18 -0.013 0.000 0.000 0.000 0.000 -0.042 -0.023 -0.024 0.066 0.000 0.067 0.037 -0.109 0.028 -0.056 0.114 -0.033
CCIKM18 0.084 0.000 0.000 0.000 0.000 0.133 0.084 -0.250 0.032 0.000 -0.066 0.040 -0.141 0.037 0.204 -0.021 -0.119
FinHC AvgF1 CHeadlines 0.503 0.500 0.500 0.500 0.500 0.695 0.500 0.769 0.818 0.504 0.607 0.716 0.626 0.690 0.526 0.952 0.957
FinQA EM CEnQA 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.008
Accuracy ConFinQA 0.068 0.004 0.000 0.008 0.000 0.021 0.000 0.000 0.021 0.000 0.008 0.030 0.276 0.013 0.046 0.169 0.228
EN DTE EN-CLS FinSA Accuracy FPB 0.388 0.231 0.000 0.286 0.001 0.741 0.696 0.780* 0.760* 0.319 0.278 0.808 0.435 0.353 0.384 0.398 0.867
F1 FPB 0.231 0.171 0.000 0.127 0.002 0.744 0.702 0.780* 0.780* 0.353 0.292 0.809 0.436 0.349 0.331 0.415 0.866
FiQA-SA 0.742 0.327 0.000 0.370 0.000 0.551 0.754 0.600* 0.800* 0.349 0.096 0.532 0.477 0.477 0.503 0.480 0.848
FinHC AvgF1 Headlines 0.600 0.600 0.600 0.600 0.600 0.655 0.600 0.770* 0.860* 0.600 0.551 0.600 0.600 0.682 0.668 0.780 0.965
FinCC F1 German 0.649 0.525 0.525 0.525 0.525 0.525 0.476 0.200* 0.550* 0.525 0.525 0.548 0.192 0.525 0.525 0.525 0.318
Australian 0.412 0.260 0.260 0.260 0.358 0.273 0.502 0.410* 0.740* 0.260 0.288 0.260 0.599 0.260 0.260 0.260 0.421
MCC German 0.231 0.000 0.000 0.000 0.000 0.000 -0.166 -0.100* -0.020* 0.000 0.000 0.012 0.002 0.000 0.000 0.000 -0.306
Australian 0.000 0.000 0.000 0.000 0.153 -0.017 0.065 0.000* 0.470* 0.000 -0.097 0.000 0.232 0.000 0.000 0.000 -0.030
EN-PRE FinSP Accuracy ACL18 0.503 0.500 0.487 0.508 0.488 0.511 0.490 0.500* 0.520* 0.510 0.483 0.492 0.493 0.515 0.519 0.525 0.502
BigData18 0.550 0.525 0.448 0.484 0.448 0.4769 0.508 0.530* 0.540* 0.469 0.452 0.472 0.496 0.555 0.520 0.483 0.450
CIKM18 0.460 0.478 0.419 0.496 0.418 0.499 0.470 0.550* 0.570* 0.469 0.421 0.430 0.461 0.563 0.556 0.534 0.493
MCC ACL18 -0.034 -0.010 0.000 0.010 0.023 0.035 -0.038 0.005* 0.020* 0.019 -0.026 0.001 -0.014 0.017 0.029 0.045 0.023
BigData18 0.002 -0.013 0.000 -0.012 0.000 0.008 -0.017 -0.025* 0.030* -0.014 0.030 0.063 -0.002 0.042 0.030 0.017 -0.012
CIKM18 -0.053 0.009 0.000 -0.028 -0.026 0.043 -0.056 0.010* 0.020* -0.034 -0.016 -0.044 -0.038 -0.001 -0.007 0.005 0.042
EN-EXT FinER Entity F1 NER 0.000 0.000 0.000 0.012 0.000 0.248 0.012 0.770* 0.830* 0.000 0.000 0.039 0.392 0.000 0.006 0.000 0.362
EN-REA FinQA EM EnQA 0.010 0.000 0.000 0.000 0.000 0.000 0.000 0.580* 0.630* 0.000 0.000 0.004 0.001 0.000 0.000 0.001 0.001
Accuracy ConFinQA 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.600* 0.760* 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000
DOF EN-OFT FinER Entity F1 FINER-ORD 0.000 0.000 0.000 0.000 0.000 0.021 0.000 0.280* 0.770* 0.000 0.000 0.008 0.000 0.000 0.000 0.000 0.000
FinST Rouge-1 ECTSUM 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
EDTSUM 0.013 0.015 0.007 0.027 0.009 0.175 0.036 0.170* 0.200* 0.032 0.008 0.051 0.033 0.009 0.018 0.017 0.015
BertScore ECTSUM 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
EDTSUM 0.522 0.447 0.445 0.483 0.445 0.599 0.536 0.660* 0.670* 0.478 0.508 0.552 0.467 0.476 0.536 0.505 0.500
BartScore ECTSUM -5.178 -5.178 -5.178 -5.178 -5.178 -5.178 -5.178 -5.178* -5.178* -5.178 -5.178 -5.178 -5.178 -5.178 -5.178 -5.178 -5.178
EDTSUM -6.996 -6.923 -7.036 -6.902 -6.957 -4.048 -6.845 -3.640* -3.620* -6.973 -7.083 -6.951 -6.964 -7.107 -6.945 -6.925 -6.927
FinDC Accuracy FOMC 0.288 0.256 0.000 0.319 0.000 0.464 0.454 0.600* 0.690* 0.252 0.208 0.557 0.476 0.385 0.373 0.488 0.429
F1 FOMC 0.207 0.172 0.000 0.323 0.000 0.470 0.441 0.640* 0.710* 0.131 0.155 0.541 0.480 0.375 0.372 0.455 0.424
Best Result Count Including ChatGPT 2 0 1 0 0 1 0 1 17 1 1 1 1 1 10 6 20
Excluding ChatGPT 2 0 1 0 0 5 0 - - 1 2 5 4 2 10 7 24
Table 10: The detailed evaluation results of different 6B-7B LLMs, baseline FinLLMs and variants of ICE-INTERN on ICE-FLARE. All results via our evaluations are the average of three runs. Results with "*" of ChatGPT-3.5(ChatGPT) and ChatGPT-4(GPT4) on English dataset are cited from  (Xie et al., 2023b). Results in bold indicate the best results across all models. Results with "__" indicate the best results among models excluding ChatGPT.
G Examples of Query Response among Various LLMs
Table 11 presents the examples of query input and response among various LLMs on some specific financial task.
Specific Task Query Input (Including Prompt Description) Model Response
Baichuan-7B ChatGPT-4 ICE-INTERN-full-7B
FinQA
浦东建设公告,公司子公司上海市浦东新区建设(集团)有限公司、上海浦兴路桥建设工程有限公司中标多项重大工程项目,中标金额总计为10.94亿元。请问中标事件对应的主体有哪些?
浦东建设、浦兴路桥、浦东建设(集团)
中标事件对应的主体有上海市浦东新区建设(集团)有限公司
上海市浦东新区建设(集团)有限公司,上海浦兴路桥建设工程有限公司
FinTS
市场普遍预期日本央行将维持其超宽松货币政策不变,但知情人士表示,周二将公布的经济展望报告中,可能上调4月起财政年度的通胀预估。能源和原材料成本的上涨已将批发价格推高至40年来的最高水平,企业已开始将上涨的成本转嫁给消费者。近期通胀温和上升,主要是投入成本上升和日元疲软推动的,而不是强劲的国内需求。日本央行行长黑田东彦曾表示,物价上涨应该伴随着工资的强劲增长。请对上述金融报道进行总结,给出具有中心主旨的摘要,回答要求尽量简短。
日本央行行长黑田东彦曾表示,物价上涨应该伴随着工资的强劲增长。
日本央行预计将上调4月起的通胀预估,尽管市场普遍预期货币政策将保持超宽松。成本上涨推高批发价格,企业已将涨成本转嫁给消费者,但通胀主要由投入成本上升和日元疲软推动,而非国内需求强劲。
知情人士:日本央行可能上调4月起财政年度的通胀预估
FinRE
新希望集团销售额约为上市公司<N>倍有余,由于背靠农业帝国“希望系”, 资本市场一直预期集团能够向新希望股份注入农业资产, 如果集团资产整合启动, 新希望股份的股。请分析上述涉及实体对(新希望股份,新希望集团)的关系。
被收购
被注资
自己
FinED
大智慧(601519)信披违规遭调查“互联网+”炒作或终结金鹰等多家基金公司高管离职 基民担忧业绩波动。请分析上述金融报道中的事件类型及及其事件主体,并以‘事件类型,事件主体’的格式输出。
套现约1.5亿元
交易违规, 大智慧
信批违规, 大智慧
FinER
本公司非执行董事孙月英女士不再担任中国远洋海运集团有限公司总会计师和中远财务有限责任公司董事长。请给出上述金融报道中存在的个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体,回答应遵循的格式’实体名称, 实体类型’。
中国远洋海运集团有限公司,ORG
中远海运发展股份有限公司, ORG
孙月英, PER; 中国远洋海运集团有限公司, ORG; 中远财务有限责任公司, ORG
FinNC
WTI原油涨幅回升至0.5%,现报75.58美元/桶。请对这个金融新闻报道进行分类,具体属于[’中国’,’外国’,’国际’,’公司’,’行业’,’大盘’,’经济’,’政策’,’政治’,’期货’,’债券’,’房地产’,’外汇’,’虚拟货币’,’新冠’,’能源’]中的哪些类别?。
输出与内容无关
国际,能源
国际期货
Table 11: Examples of of query input and response among various LLMs on some specific financial task of Question Answering(FinQA), Text Summarization(FinTS), Relationship Extraction(FinRE), Event Detection(FinED), Entity Recognition(FinER) and News Classification(FinNC).